OpenStudy (anonymous):

find the 6 six trig functions of (2√2/3),(1/3)

OpenStudy (jdoe0001):

$$\bf \left(\frac{2\sqrt{2}}{3}, \frac{1}{3}\right) \implies (x, y) \implies (adjacent, opposite) $$ so use the trig identities :) http://faculty.bucks.edu/farberb/trig_i1.gif